Find the probability and interpret the results. If​ convenient, use technology to find the probability.
The population mean annual salary for environmental compliance specialists is about ​$62 comma 000
.
A random sample of 34

specialists is drawn from this population. What is the probability that the mean salary of the sample is less than ​$59 comma 500
​?
Assume sigma
equals​$6,200
.
The probability that the mean salary of the sample is less than ​$59, 500


Answer by stanbon(75887)

Find the probability and interpret the results. If​ convenient, use technology to find the probability.
The population mean annual salary for environmental compliance specialists is about ​$62,000.
A random sample of 34 specialists is drawn from this population.
What is the probability that the mean salary of the sample is less than ​$59,500​?
Assume sigma equals ​$6,200.
z(59,500) = (59500-62000)/[6200/sqrt(34)] = -2.3512
Ans: P(x-bar < 59500) = P(z < -2.3512) = normalcdf(-100,-2.3512) = 0.0094
-----
Cheers,
Stan H.
----------